Compute the z-statistic for each given claim (p), the observed proportion (^p), and the given

sample size (𝑛).


𝑝 = 0.2, ^p= 0.18, 𝑛 = 50



1
Expert's answer
2022-05-30T09:29:31-0400

"z=\frac{\^p-p}{\sqrt{\frac{p(1-p)}{n}}}"

"z=\frac{0.18-0.2}{\sqrt{\frac{0.2(1-0.2)}{50}}}=-0.353"